A company produced 300 appliances in the first week of the month. Because it received additional machinery, its production increased 50% from the first week to the second week. How many appliances did the company produce the second week ?

Answer by alicealc(293)

the increase = 50% * 300 = 50/100 * 300 = 0.5 * 300 = 150
so in the second week, the company can produce 300 + 150 = 450 appliances